# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v2

@@ 6.1 Simulation Designs
    $M=A(\mathbf{B}_{1}X+U^{\prime})+(1-A)(\mathbf{B}_{0}X+U)\,,$
@@
    $Y=A(X^{\top}\beta_{1}+M^{\top}\gamma_{1}+{\epsilon})+(1-A)(X^{\top}\beta_{0}+M^{\top}\gamma_{0}+{\epsilon}^{\prime})\,,$
@@
The number of samples $n$ varies over $\{250,500,750,1000,1250\}$. For each value of $n$ in this set, we vary $p=q$ over the set $\{50,400,750,1250\}$. Furthermore, the noise variance $\sigma^{2}$ is varied over the set $\{0.1,0.5,1\}$. Recall that our parameter of interest is $\theta_{0}=\bar{X}^{\top}(\beta+\mathbf{B}_{0}^{\top}\gamma_{1})$ where $\bar{X}$ is the sample average of the covariates. We keep $X$ and other parameters (e.g., $(\beta_{j},\gamma_{j},\alpha,\mathbf{B}_{j})$ for $j\in\{0,1\}$) fixed in the simulation study to keep the parameter of interest remains unchanged.
We follow Algorithm 1 to compute our estimator. The Lasso estimates in Step 1 and Step 3 are obtained using the R package glmnet, where the tuning parameter is selected by cross-validation using the lambda.1se rule from R function cv.glmnet (Friedman et al., (2010)). The debiasing weights in Step 2 and Step 4 by solving the corresponding optimization problem via R package pogs with $K_{1}=K_{2}=2.75$. This particular choice of $(K_{1},K_{2})$ is motivated by our sensitivity analysis, as elaborated in Table 4 - 7. This choice remains consistent with previous studies, e.g., Javanmard and Montanari, (2014).
Figure 2: Histogram and QQ plot of $\sqrt{n}(\hat{\theta}-\theta)$ under $n=1000,p+q=800,\sigma_{\epsilon}=0.5$.
The simulation results are summarized in Tables 1. Table 1 presents the root-mean-squared error (RMSE) and standard deviation (SD) of both the naive estimator and our proposed debiased estimator (computed using Algorithm 1), averaged over 400 Monte Carlo replications, and reports the corresponding standard deviations of the estimators across these replications. The results indicate that the debiased estimator consistently achieves a lower RMSE than the naive estimator, with the performance gap becoming increasingly prominent as the sample size as $n$ grows. Furthermore, the standard deviations of the two estimators are comparable (as evident from Table 1), suggesting that the improved RMSE of our debiased estimator is attributable to the successful bias correction. To further illustrate the performance of the estimators, we display histograms of the naive (red) and debiased (blue) estimators in the left column of Figure 2, based on 400 Monte Carlo replications. The estimators are centered around the true parameter and scaled by $\sqrt{n}$, where $n$ is the number of samples. The histogram clearly reveals that the naive estimator is substantially biased, while the debiased estimator is centered around zero as a result of successful bias correction. Additionally, the Q-Q plot of the debiased estimator, shown in the right column of Figure 2, aligns closely with the standard normal distribution, providing empirical evidence for its asymptotic normality and supporting our theoretical results.
@@ 6.2 Real Data Experiments
  • We selected $20$ clinically relevant features, $X$ (e.g., age, sex, etc.) for our study by excluding all those features which has more than $100$ missing values. See the Supplementary Materials for the complete list of the chosen features. We then imputed the remaining missing values; for numerical features (11), we used softImpute (Mazumder et al., , 2010) and for categorical variables (9), we used mice (Van Buuren and Groothuis-Oudshoorn, , 2011) imputation. Furthermore, we considered second, third-order and fourth-order polynomial features and interaction terms between categorical and numerical variables for better expressiveness. These operations finally lead to a $449$ dimensional covariate.
@@
After the data pre-processing, we are left with $n=738$ units, each with $449$ features $(X)$ and $124,333$ mediators ($M$). The treatment indicator ($A$) is the smoking status of the units ($A=1$ indicates a smoker). Among these samples, there were $42.6\%(315)$ are females and $57.3\%(423)$ are males. The patients’ ages range from 33 to 90 years, with a median age of 67 years. The median survival time is 603 days (1.65 years).
@@
[…]
@@
We compared the results with a fixed-dimensional OLS-based approach. Note that if both $X$ and $M$ are low-dimensional in Equation (6.2), we can easily estimate the NDE, NIE (and consequently the total effect) by computing ordinary least squared (OLS) esimators via regressing i) $Y$ on $(X,M)$ (on treatment and control observations separately), and ii) $M$ on $X$ (on treatment and control observations separately). To make the problem finite-dimensional, we selected CpG sites (cg19757631, cg05147638, cg24720672, cg08108679, cg05575921, cg24859433) which were identified as significant mediators in previous studies (Cui et al., , 2021; Zhang et al., 2021a, ; Zhang et al., 2021b, ), as our $M$, and $20$ clinically relevant features as our $X$ (i.e., we do not consider their higher-order terms and interactions).
In Table 3, we present the results in a fixed-dimensional setting. The results reveal that the OLS-based approach produces conclusions about the indirect and total effects that are contrary to intuition. The estimate of the indirect and total effect imply that smokers have $128.80\%$ ($e^{0.2531}\times 100\%$) (mediated through DNA methylation), and $126.52\%$ ($e^{0.2353}\times 100\%$) of survival time of non-smokers, i.e., smoking increases $\approx 29\%$ $=100\times(1-e^{0.2531})\%$ (corresponds to the NIE) and $\approx 27\%$ $=100\times(1-e^{0.2353})\%$ of the survival time of lung cancer patients. Furthermore, the confidence interval for the direct effect and total effect includes zero, suggesting a lack of statistical significance. These findings further emphasize the limitations of low-dimensional mediator selection and conventional modeling approaches, reinforcing the importance of incorporating high-dimensional methodologies for accurate mediation analysis in complex biological systems.
@@ Appendix A Proofs
We begin our analysis by using the contrast $a=\begin{bmatrix}\bar{X}&(\mathbf{\hat{B}}_{0}\bar{X})\end{bmatrix}$ and $b=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, define
$\hat{\vartheta}$ $=a^{\top}\hat{\phi}_{1}+\tau_{1}^{\top}(Y_{t}-\mathbf{X}_{t}^{\top}\hat{\beta}_{1}-\mathbf{M}_{t}^{\top}\hat{\gamma}_{1})+\bar{X}^{\top}\hat{b}+\tau_{2}^{\top}(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}\hat{b})$
$=a^{\top}\hat{\phi}_{1}+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}\cdot\hat{\phi}_{1})+\bar{X}^{\top}\hat{b}+\tau_{2}^{\top}(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}\hat{b})\,.$
@@
$a^{\top}\hat{\phi}_{1}+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}\cdot\hat{\phi}_{1})$
$=a^{\top}\phi_{1}+a^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}\cdot{\phi}_{1})+\tau_{1}^{\top}\mathbf{W}_{t}\left(\phi_{1}-\hat{\phi}_{1}\right)$
@@
$\bar{X}^{\top}\hat{b}+\tau_{2}^{\top}\left(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}\hat{b}\right)$
$=\bar{X}^{\top}b+\bar{X}^{\top}(\hat{b}-b)+\tau_{2}^{\top}\left(\mathbf{M}_{c}^{\top}\hat{\gamma}_{1}-\mathbf{X}_{c}^{\top}b+\mathbf{X}_{c}^{\top}(b-\hat{b})\right)$
$=\bar{X}^{\top}b+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}$
$=\bar{X}^{\top}\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,.$
@@
$\hat{\vartheta}=\bar{X}^{\top}\beta_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\gamma_{1}+\bar{X}^{\top}\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,.$
@@
$=\bar{X}^{\top}\beta_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(\mathbf{B}_{0}\bar{X}\right)^{\top}\hat{\gamma}_{1}+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}$
$=\bar{X}^{\top}\beta_{1}+\left(\mathbf{B}_{0}\bar{X}\right)^{\top}\gamma_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\hat{\gamma}_{1}-\gamma_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)$
$~{}~{}~{}~{}~{}~{}~{}+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}$
$=\theta_{0}+\left(\hat{\mathbf{B}}_{0}\bar{X}-\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,.$
@@
$\hat{\theta}_{0}-\theta_{0}=\left(\hat{\mathbf{B}}_{0}\bar{X}-\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}U^{\top}\hat{\gamma}_{1}\,,$
@@
$\tau_{1,i}^{*}=\frac{1}{n_{t}}\check{a}^{\top}\Sigma_{t}^{-1}(\mathbf{W}_{t,i}-\mu_{t}),\qquad\tau_{2,i}^{*}=\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1}(\mathbf{X}_{c,i}-\mu_{c}),\qquad 1\leq i\leq n\,.$ (A.1)
where $\check{a}=[\mu\quad\hat{\mathbf{B}}_{0}\mu],\mu=\mathbb{E}[X]$. We start with $\tau_{1}^{*}$ and show that with high probability, it satisfies the constraints of Equation (5.4), i.e.
@@
$\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}$ $\leq\|a-\check{a}\|_{\infty}+\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\,.$ (A.3)
@@
$\|a-\check{a}\|_{\infty}=\|[\bar{X}\quad\hat{\mathbf{B}}_{0}\bar{X}]-[\mu\quad\hat{\mathbf{B}}_{0}\mu]\|_{\infty}=\max\{\|\bar{X}-\mu\|_{\infty},\|\hat{\mathbf{B}}_{0}(\bar{X}-\mu)\|_{\infty}\}\,.$ (A.4)
As we have assumed $X_{i}$’s sub-Gaussian (Assumption 5.6), a simple application of Hoeffding’s inequality along with a union bound yields $\|\bar{X}-\mu\|_{\infty}\leq K_{2}\sqrt{\log{(p)}/n}$ with probability approaching to 1. For the other term, we use the fact that with probability going to 1, $\|\hat{\mathbf{B}}_{0}-\mathbf{B}_{0}\|_{1,\infty}\leq C_{1}s\sqrt{\log(p)/n_{c}}$ for some constant $C_{1}>0$, which is later established in Lemma A.1. Using this information, we bound the $\|\hat{\mathbf{B}}_{0}(\bar{X}-\mu)\|_{\infty}$ as follows:
$\|\hat{\mathbf{B}}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\mu\|_{\infty}$ $=\|\mathbf{B}_{0}\bar{X}-\mathbf{B}_{0}\mu-\hat{\mathbf{B}}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\mu-\mathbf{B}_{0}\bar{X}+\mathbf{B}_{0}\mu\|_{\infty}$
$\leq\|\mathbf{B}_{0}(\bar{X}-\mu)\|_{\infty}+\|\hat{\mathbf{B}}_{0}-\mathbf{B}_{0}\|_{1,\infty}\|\bar{X}-\mu\|_{\infty}$
$=\max_{1\leq j\leq q}\left|\left\langle\mathbf{B}_{0,j\cdot},\bar{X}-\mu\right\rangle\right|+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq\max_{1\leq j\leq q}\|\mathbf{B}_{0,j\cdot}\|_{1}\|\bar{X}-\mu\|_{\infty}+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq 2\|\mathbf{B}_{0}\|_{1,\infty}\|\bar{X}-\mu\|_{\infty}+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq C_{B}K_{2}\sqrt{\frac{\log{(p})}{n}}+K_{2}C_{1}s\frac{\log{(p)}}{n_{c}}:=K_{3}\sqrt{\frac{\log{(p})}{n_{c}}}\,,$
@@
$\|a-\check{a}\|_{\infty}\leq(K_{2}\vee K_{3})\sqrt{\frac{\log{(p})}{n_{c}}}\,.$ (A.5)
Next, we provide an upper bound on $\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}$ in Equation (A.4). Towards that end, we write $(\mathbf{W}_{t}^{\top}\tau_{1}^{*})$ as:
$\mathbf{W}_{t}^{\top}\tau_{1}^{*}$ =$=\sum_{\{i:A_{i}=1\}}\mathbf{W}_{t,i}\tau_{1,i}^{*}=\frac{1}{n_{t}}\sum_{i}\mathbf{W}_{t,i}(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}$
$=\underbrace{\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(\mathbf{W}_{t,i}-\mu_{t})(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}}_{:=\xi_{1}}+\underbrace{\mu_{t}\frac{1}{n_{t}}\sum_{i}(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}}_{:=\xi_{2}}\,.$ (A.6)

### v3

@@ 6.1 Simulation Designs
    $M=(1-A)(\mathbf{B}_{0}X+U)+A(\mathbf{B}_{1}X+U^{\prime})\,$,
@@
    $Y=(1-A)(X^{\top}\beta_{0}+M^{\top}\gamma_{0}+\epsilon^{\prime})+A(X^{\top}\beta_{1}+M^{\top}\gamma_{1}+\epsilon)\,$,
@@
The number of samples $n$ varies over $\{250,500,750,1000,1250\}$. For each value of $n$ in this set, we vary $p=q$ over the set $\{50,400,750,1250\}$. Furthermore, the noise variance $\sigma^{2}$ is varied over the set $\{0.1,0.5,1\}$. Recall that our parameter of interest is $\theta_{0}=\bar{X}^{\top}(\beta+\mathbf{B}_{0}^{\top}\gamma_{1})$ where $\bar{X}$ is the sample average of the covariates. We keep $X$ and other parameters (e.g., $(\beta_{j},\gamma_{j},\alpha,\mathbf{B}_{j})$ for $j\in\{0,1\}$) fixed in the simulation study to keep the parameter of interest unchanged.
@@
We follow Algorithm 1 to compute our estimator. The Lasso estimates in Step 1 and Step 3 are obtained using the R package glmnet, where the tuning parameter is selected by cross-validation using the lambda.1se rule from R function cv.glmnet (Friedman et al., (2010)). The debiasing weights in Step 2 and Step 4 are solved by solving the corresponding optimization problem via the R package pogs with $K_{1}=K_{2}=2.75$. This particular choice of $(K_{1},K_{2})$ is motivated by our sensitivity analysis, as elaborated in Table 4 - 7 in the supplementary material. This choice remains consistent with previous studies, e.g., Javanmard and Montanari, (2014).
Figure 2: Histogram and $Q$ plot of $\sqrt{n}(\hat{\theta}-\theta)$ under $n=1000,p+q=800,\sigma_{\epsilon}=0.5$.
The simulation results are summarized in Tables 1. This table presents the root-mean-squared error (RMSE) and standard deviation (SD) of both the naive estimator and our proposed debiased estimator (computed using Algorithm 1), averaged over 400 Monte Carlo replications, and reports the corresponding standard deviations of the estimators across these replications. The results indicate that the debiased estimator consistently achieves a lower RMSE than the naive estimator, with the performance gap becoming increasingly prominent as the sample size $n$ grows. Furthermore, the standard deviations of the two estimators are comparable (as evident from Table 1), suggesting that the improved RMSE of our debiased estimator is attributable to the successful bias correction. To further illustrate the performance of the estimators, we display histograms of the naive (red) and debiased (blue) estimators in the left column of Figure 2, based on 400 Monte Carlo replications. The estimators are centered around the true parameter and scaled by $\sqrt{n}$, where $n$ is the number of samples. The histogram clearly reveals that the naive estimator is substantially biased, while the debiased estimator is centered around zero as a result of successful bias correction. Additionally, the Q-Q plot of the debiased estimator, shown in the right column of Figure 2, aligns closely with the standard normal distribution, providing empirical evidence for its asymptotic normality and supporting our theoretical results.
@@ 6.2 Real Data Experiments
  • We selected $20$ clinically relevant features, $X$ (e.g., age, sex, etc.; see full list in Supplementary Material Section 8) for our study by excluding all those features which has more than $100$ missing values. See the Supplementary Materials for the complete list of the chosen features. We then imputed the remaining missing values; for numerical features (11), we used softImpute (Mazumder et al., , 2010) and for categorical variables (9), we used mice (Van Buuren and Groothuis-Oudshoorn, , 2011) imputation. Furthermore, we considered second, third-order and fourth-order polynomial features and interaction terms between categorical and numerical variables for better expressiveness. These operations finally lead to a $449$ dimensional covariate.
@@
After the data pre-processing, we are left with $n=738$ units, each with $449$ features $(X)$ and $124,333$ mediators ($M$). The treatment indicator ($A$) is the smoking status of the units ($A=1$ indicates a smoker). Among these samples, there are $42.6\%(315)$ females and $57.3\%(423)$ males. The patients’ ages range from 33 to 90 years, with a median age of 67 years. The median survival time is 603 days (1.65 years).
@@
We then apply Algorithm 1 to estimate $\mathbb{E}[Y^{(1,0)}],\mathbb{E}[Y^{(1,1)}]$ and $\mathbb{E}[Y^{(0,0)}]$ (see supplementary material B for more details). Table 2 summarizes the estimates of NDE, NIE, and the total effect of smoking on (log) survival time. The estimate of the total effect implies that, on average, smokers have 66.70% ($e^{-0.4049}\times 100\%$) of the survival time of non-smokers, i.e., smoking reduces $\approx\textbf{33\%}$ of the survival time of lung cancer patients. Out of this total effect, our results demonstrated that $\approx\textbf{28\%}=100\times(1-e^{-0.3319})\%$ corresponds to the indirect effect (NIE), i.e., the impact of smoking on survival time, mediated via DNA methylation. The remaining $5\%$ (NDE) can be attributed to the factors beyond the methylation mechanism considered in this analysis. Since a large proportion of the total effect is driven by the indirect effect, our findings suggest that smoking primarily affects survival time through DNA methylation rather than direct physiological pathways. This highlights the crucial role of DNA methylation as a mediator in the relationship between smoking and survival.
@@
We compared the results with a fixed-dimensional OLS-based approach. Note that if both $X$ and $M$ are low-dimensional in Equation (6.2), we can easily estimate the NDE, NIE (and consequently the total effect) by computing ordinary least squared (OLS) estimators via regressing i) $Y$ on $(X,M)$ (on treatment and control observations separately), and ii) $M$ on $X$ (on treatment and control observations separately). To make the problem finite-dimensional, we selected CpG sites (cg19757631, cg05147638, cg24720672, cg08108679, cg05575921, cg24859433) which were identified as significant mediators in previous studies (Cui et al., , 2021; Zhang et al., 2021a, ; Zhang et al., 2021b, ), as our $M$, and $20$ clinically relevant features as our $X$ (i.e., we do not consider their higher-order terms and interactions).
In Table 3, we present the results in a fixed-dimensional setting. The results reveal that the OLS-based approach produces conclusions about the indirect and total effects that are contrary to intuition. The estimate of the indirect and total effect imply that smokers have $128.80\%$ ($e^{0.2531}\times 100\%$) (mediated through DNA methylation), and $126.52\%$ ($e^{0.2353}\times 100\%$) of survival time of non-smokers, i.e., smoking increases $\approx 29\%$ $100\times(1-e^{0.2531})\%$ (corresponds to the NIE) and $\approx 27\%$ $100\times(1-e^{0.2353})\%$ of the survival time of lung cancer patients. Furthermore, the confidence interval for the direct effect and total effect includes zero, suggesting a lack of statistical significance. These findings further emphasize the limitations of low-dimensional mediator selection and conventional modeling approaches, reinforcing the importance of incorporating high-dimensional methodologies for accurate mediation analysis in complex biological systems.
@@ Appendix A Proofs
We begin our analysis by using the contrast $a=(\bar{X}^{\top}\quad(\mathbf{\hat{B}}_{0}\bar{X})^{\top})$ and $b=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, define
[…]
@@
$a^{\top}\hat{\phi}_{1}+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}\hat{\phi}_{1})$
$=a^{\top}\phi_{1}+a^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\tau_{1}^{\top}(Y_{t}-\mathbf{W}_{t}{\phi}_{1})+\tau_{1}^{\top}\mathbf{W}_{t}\left(\phi_{1}-\hat{\phi}_{1}\right)$
@@
[…]
$=\bar{X}^{\top}b+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{2}^{\top}\bm{U}\hat{\gamma}_{1}$
$=\bar{X}^{\top}\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)+\tau_{2}^{\top}\bm{U}\hat{\gamma}_{1}\,.$
@@
$\hat{\vartheta}$ $=\bar{X}^{\top}\beta_{1}+\left(\hat{\mathbf{B}}_{0}\bar{X}\right)^{\top}\gamma_{1}+\bar{X}^{\top}\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}(\hat{b}-b)$
$~~~~~~+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}\bm{U}\hat{\gamma}_{1}\,.$
@@
$=\bar{X}^{\top}\beta_{1}+(\hat{\mathbf{B}}_{0}\bar{X})^{\top}(\gamma_{1}-\hat{\gamma}_{1})+(\mathbf{B}_{0}\bar{X})^{\top}\hat{\gamma}_{1}$
$\quad+(a-\mathbf{W}_{t}^{\top}\tau_{1})^{\top}(\hat{\phi}_{1}-\phi_{1})+(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2})^{\top}(\hat{b}-b)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}\bm{U}\hat{\gamma}_{1}$
$=\theta_{0}+(\hat{\mathbf{B}}_{0}\bar{X}-\mathbf{B}_{0}\bar{X})^{\top}(\gamma_{1}-\hat{\gamma}_{1})+(a-\mathbf{W}_{t}^{\top}\tau_{1})^{\top}(\hat{\phi}_{1}-\phi_{1})$
$\quad+(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2})^{\top}(\hat{b}-b)+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}\bm{U}\hat{\gamma}_{1}\,.$
@@
$\hat{\theta}_{0}-\theta_{0}$ $=\left(\hat{\mathbf{B}}_{0}\bar{X}-\mathbf{B}_{0}\bar{X}\right)^{\top}\left(\gamma_{1}-\hat{\gamma}_{1}\right)+\left(a-\mathbf{W}_{t}^{\top}\tau_{1}\right)^{\top}\left(\hat{\phi}_{1}-\phi_{1}\right)+\left(\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\right)^{\top}\left(\hat{b}-b\right)$
$\quad+\tau_{1}^{\top}\epsilon+\tau_{2}^{\top}\mathbf{U}\hat{\gamma}_{1}\,$,
@@
$\tau_{1,i}^{*}=\frac{1}{n_{t}}\tilde{a}^{\top}\Sigma_{W,t}^{-1}(W_{t,i}-\mu_{W,t}),\qquad\tau_{2,i}^{*}=\frac{1}{n_{c}}\mu_{X}^{\top}\Sigma_{X,c}^{-1}(X_{c,i}-\mu_{X,c}),\qquad 1\leq i\leq n\,.$ (A.1)
where $\tilde{a}=(\mu_{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\mu_{X})^{\top})^{\top}$, and $\mu_{X}=\mathbb{E}[X]$. We start with $\tau_{1}^{*}$ and show that with high probability, it satisfies the constraints of Equation (5.4), i.e.
@@
$\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}$ $\leq\|a-\tilde{a}\|_{\infty}+\|\tilde{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\,.$ (A.3)
@@
$\|a-\tilde{a}\|_{\infty}$ $=\|(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}-(\mu_{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\mu_{X})^{\top})^{\top}\|_{\infty}$ (A.4)
$=\max\{\|\bar{X}-\mu_{X}\|_{\infty},\|\hat{\mathbf{B}}_{0}(\bar{X}-\mu_{X})\|_{\infty}\}\,.$
As we have assumed $X_{i}$’s sub-Gaussian (Assumption 5.7), a simple application of Hoeffding’s inequality along with a union bound yields $\|\bar{X}-\mu_{X}\|_{\infty}\leq K_{2}^{\prime}\sqrt{\log{(p)}/n}$ with probability approaching to 1. For the other term, we use the fact that with probability going to 1, $\|\hat{\mathbf{B}}_{0}-\mathbf{B}_{0}\|_{1,\infty}\leq C_{1}s\sqrt{\log(p)/n_{c}}$ for some constant $C_{1}>0$, which is later established in Lemma A.1. Using this information, we bound the $\|\hat{\mathbf{B}}_{0}(\bar{X}-\mu_{X})\|_{\infty}$ as follows:
$\|\hat{\mathbf{B}}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\mu_{X}\|_{\infty}$ $=\|\mathbf{B}_{0}\bar{X}-\mathbf{B}_{0}\mu_{X}-\hat{\mathbf{B}}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\mu_{X}-\mathbf{B}_{0}\bar{X}+\mathbf{B}_{0}\mu_{X}\|_{\infty}$
$\leq\|\mathbf{B}_{0}(\bar{X}-\mu_{X})\|_{\infty}+\|\hat{\mathbf{B}}_{0}-\mathbf{B}_{0}\|_{1,\infty}\|\bar{X}-\mu_{X}\|_{\infty}$
$=\max_{1\leq j\leq q}\left|\left\langle B_{0,j\cdot},\bar{X}-\mu_{X}\right\rangle\right|+K_{2}^{\prime}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq\max_{1\leq j\leq q}\|B_{0,j\cdot}\|_{1}\|\bar{X}-\mu_{X}\|_{\infty}+K_{2}^{\prime}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq 2\|\mathbf{B}_{0}\|_{1,\infty}\|\bar{X}-\mu_{X}\|_{\infty}+K_{2}^{\prime}C_{1}s\frac{\log{(p)}}{n_{c}}$
$\leq C_{B}K_{2}^{\prime}\sqrt{\frac{\log{(p})}{n}}+K_{2}^{\prime}C_{1}s\frac{\log{(p)}}{n_{c}}:=K_{3}\sqrt{\frac{\log{(p})}{n_{c}}}\,$,
@@
$\|a-\tilde{a}\|_{\infty}\leq(K_{2}^{\prime}\vee K_{3})\sqrt{\frac{\log{(p})}{n_{c}}}\,.$ (A.5)
Next, we provide an upper bound on $\|\tilde{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}$ in Equation (A.4). Towards that end, we write $(\mathbf{W}_{t}^{\top}\tau_{1}^{*})$ as:
$\mathbf{W}_{t}^{\top}\tau_{1}^{*}$ $=\sum_{\{i:A_{i}=1\}}W_{t,i}\tau_{1,i}^{*}=\frac{1}{n_{t}}\sum_{i}W_{t,i}(W_{t,i}-\mu_{W,t})^{\top}\Sigma_{W,t}^{-1}\tilde{a}$
$=\underbrace{\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(W_{t,i}-\mu_{W,t})(W_{t,i}-\mu_{W,t})^{\top}\Sigma_{W,t}^{-1}\tilde{a}}_{:=\xi_{1}}+\underbrace{\mu_{W,t}\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(W_{t,i}-\mu_{W,t})^{\top}\Sigma_{W,t}^{-1}\tilde{a}}_{:=\xi_{2}}$